### v1

# Quantum Computational Insurance and Actuarial Science

Abstract
In recent years, quantum computation has been rapidly advancing, driving a technological revolution with significant potential across various sectors, particularly in finance. Despite this, the insurance industry, an essential tool for mitigating unforeseen risks and losses, has received limited attention. This paper provides an initial exploration into the realm of quantum computational insurance and actuarial science. After introducing key insurance models and challenges, we examine quantum algorithms designed to address complex insurance issues. Our study includes experimental and numerical demonstrations of quantum applications in non-life insurance, life insurance, and reinsurance. Additionally, we explore the timeline for quantum insurance, the development of quantum-enhanced insurance products, and the challenges posed by quantum computational advancements.
I Introduction
Insurance serves as a financial mechanism that helps individuals avoid potential losses and hedge against risks from specific sources [1]. It encompasses three main branches: non-life insurance, life insurance, and risk management. In the branch of non-life insurance, the central tasks are to price the premiums paid by an individual policyholder, to model the claims for an insurer to afford when an insured event happens, and to estimate the liability considering the Incurred But Not Reported (IBNR) and Reported But Not Paid (RBNP) claims [2]. While the life insurance branch pays attention to insure life against disability-and-death risk by mortality forecasting [3, 4, 5], as well as against live-too-long risk by life annuity pricing [3, 6]. In the risk management branch, reinsurance and risk indicators are common tools for transferring and quantifying the risks [7]. Ultimately, the core tasks of insurance practice and actuarial science revolve around modeling and evaluating random risky events [8]. To achieve these objectives, a variety of computational methods are employed.
Nevertheless, the traditional insurance industry and academia are encountering challenges due to the rapidly growing size of data and increasingly complex models, which will cost exponentially increased computational resources. One exciting answer to these problems is quantum computation [9], which has been widely applied in chemistry [10], biology [11, 12], materials [13], drug design [14], and especially finance [15, 16, 17]. Specifically, the quantum simulation of continuous-time stochastic processes is introduced to model the collective risk and the insurance cash flow [18, 19]. However, quantum computational insurance is still in its infancy compared with other financial branches: few quantum algorithms focus on insurance practice, and no systematic study on actuarial science has been proposed [15, 16, 17].
This article provides an early yet insightful exploration of addressing challenges in insurance practice and actuarial science with quantum computation, We respectively review the main problems in non-life insurance, life insurance, and risk management. Then quantum algorithms for these tasks are discussed. Moreover, we also discuss them with respect to the noisy intermediate-scale quantum (NISQ) era [20] and fault-tolerant quantum computation (FTQC) era [21] based on their requirement for quantum resources, offering a timetable for quantum computational insurance.
We also numerically and experimentally study several quantum insurance applications corresponding to mainstream insurance branches, providing further evidence of the feasibility of quantum computational methods for realistic insurance problems. Initially, we present numerical simulations for the excess evaluation problem. Following this, we experimentally demonstrate solutions for reinsurance type allocation and the Lee-Carter Model [4] using the superconducting quantum processor “Wukong” [22]. This marks the first implementation of an insurance-specific quantum algorithm on a quantum processor using real datasets.
Beyond the applicability and demonstrability of quantum insurance, we further explore the influence and impact of quantum technology on the insurance industry. We consider the three main developmental stages of quantum insurance and highlight how quantum computation is driving the creation of new types of insurance products. Despite these advancements, some fundamental challenges still need to be addressed.
II Basics in quantum computation
This part briefly reviews some basic concepts in quantum computation. The first is the quantum bit (qubit), the basic information in quantum computation. It is different from the classical bit as it can be in a superposition of states, a two-dimensional vector in the Hilbert space:
$|\psi\rangle=\alpha|0\rangle+\beta|1\rangle,$ (1)
where $|0\rangle$ and $|1\rangle$ are the two basis states, $\alpha$ and $\beta$ are complex amplitudes satisfying $|\alpha|^{2}+|\beta|^{2}=1$. The multi-qubit state is within the tensor product of individual qubit space, indicating that an $n$-qubit state represents a $2^{n}$-dimensional vector in the Hilbert space.
In quantum computation, we perform operations on qubits to accomplish specific tasks. The basic operations in the quantum circuit model are quantum logic gates, which are mathematically described by unitary matrices. Specifically, the single-qubit and two-qubit gates are commonly used. For instance, the single-qubit $X$ gate is the NOT gate:
$X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\to\begin{cases}X|0\rangle=&|1\rangle,\\ X|1\rangle=&|0\rangle.\end{cases}$ (2)
The two-qubit Controlled-X gate applies the $X$ gate on the target qubit according to the state of the control qubit:
$CX=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}\to\begin{cases}CX|00\rangle=&|00\rangle,\\ CX|01\rangle=&|01\rangle,\\ CX|10\rangle=&|11\rangle,\\ CX|11\rangle=&|10\rangle.\end{cases}$ (3)
An $n$-qubit quantum logic gate can process the $2n$-dimensional information in one step, outperforming classical logic gates. It can also cause entanglements or interference between qubits. These features make quantum computation potential to achieve speedups over classical computations. Quantum algorithms based on these principles have been designed for various regions including chemistry, biology, finance, etc.
III Problems related to insurance
III.1 Non-life Insurance
Non-life insurance typically covers risks associated with damage to property, legal liabilities, and medical expenses. Specifically:
Collective Risk Model (CRM)
In CRM [23], the aggregate amount of claims is given by $\sum_{i=1}^{N(t)}Y_{i}$, where $N(t)$ is the number of claims up to time $t$, modeled by a discrete count process (e.g., a Poisson process), and $Y_{i}$ is the size of the $i$-th claim, modeled by a continuous stochastic process. A primary concern within CRM is the risk of ruin, which examines the potential for an insurance company to go bankrupt. where the ruin probability is $P_{\text{ruin}}(u)=P(u(t)<0)$, with $u(t)$ the surplus of the insurance company at time $t$ and $u=u(0)$ the initial surplus. In CRM, the arrival of claims is typically assumed to be independent.
Policy Excess and Excess of Loss Reinsurance
In non-life insurance such as motor vehicle insurance, it is common to consider policies with a Policy Excess [24]. This means that the risk is divided between the insured party and the insurer according to the contract. Mathematically, if $X$ is the total claim amount and $E$ is the excess amount, the payment by the insurer $I$ can be modeled as $I=\max(X-E,0)$. A similar model is used for Excess of Loss Reinsurance, where the risk is shared between the insurer and the reinsurer. The payment by the reinsurer $R$ can be expressed as $R=\max(X-M,0)$, where $M$ is the retention limit. One common important task in these cases is to model the payments of the insured party, the insurer, and the reinsurer.
Generalized Linear Model
The generalized linear model [25] is widely used for short-term insurance pricing. The expectation of the outcome $Y$ can be derived as:
$\mathbb{E}[Y|X]=\mu=g^{-1}(X\beta),$ (4)
where $g$ is the link function and $X\beta$ is the linear predictor. Different link functions and corresponding models are used for different insurance situations. For example, Poisson regression is typically used for modeling the frequency of claims, while Gamma regression may be applied to model the severity of claims.
IBNR
It often happens that some insurance policies have incurred claims that have not yet been reported [26]. The Chain-Ladder Model [27] is a mainstream model used to estimate the IBNR amount. In this model, the claims amount $A_{i,j}$ is aggregated by accident year $i$ and payment year $j$. The aggregated claims amounts are assumed to be linked by a chain-ladder factor $f_{j}$ as $A_{i,j+1}\sim f_{j}A_{i,j}$ and the core problem is to estimate the chain-ladder factors $f_{j}$.
Other related Problems
In addition to the previously discussed issues, non-life insurance also deals with various other problems. These include catastrophe modeling [28], which involves estimating the impact of large-scale events such as natural disasters; premium calculation, which involves determining the appropriate premium to charge for a given level of risk; reserve setting [29], which entails estimating the reserves needed to cover future claims; and solvency assessment [30], which evaluates the financial health and solvency of the insurance company. Addressing these challenges often requires the application of advanced statistical and actuarial techniques, alongside robust computational methods.
III.2 Life Insurance
Life insurance is a financial product that provides a payout to beneficiaries upon the death of the insured individual. It serves as a crucial tool for financial planning, offering security and peace of mind to policyholders and their families. The valuation of life insurance contracts is a complex and computationally intensive task, involving various factors such as mortality rates, lapse rates, and financial returns.
Valuation of Insurance Contracts
A key challenge in life insurance is accurately valuing insurance contracts [31]. The expected cash outflow of a whole life insurance policy can be modeled as
$\text{E}(C)=\sum_{t=1}^{T}q(t)\cdot P_{\text{lapse}}(t),$ (5)
where $\text{E}(C)$ is the expected cash outflow, $q(t)$ and $P_{\text{lapse}}(t)$ are the payment and the probability of lapse at time $t$. Since policyholders sometimes cancel their contracts, the dynamic lapse rate $\lambda(t)$ is considered:
$P_{\text{lapse}}(t)=1-e^{-\lambda(t)t}.$ (6)
These equations help capture the complexities involved in contract valuation by integrating both the timing of payments and the likelihood of policy termination over time.
Mortality Forecasting
Multivariate statistical analysis and data science techniques are introduced to solve the mortality forecasting problems, which have many variants. For instance, the Lee-Carter model [4] is a popular approach for mortality forecasting:
$\ln m_{x,t}=\alpha_{x}+\beta_{x}\kappa_{t}+\epsilon_{x,t},$ (7)
where $m_{x,t}$ is the mortality rate for age $x$ at time $t$, $\alpha_{x}=(\sum_{x=1}^{T}m_{x,t})/T$ represents the age-specific component, $\beta_{x}$ captures the sensitivity to the time-varying index $\kappa_{t}$, and $\epsilon_{x,t}$ is the error term. This model and its variants are extensively discussed in [32].
Multi-period Optimization
For life insurance, especially for life annuities, Multi-period Optimization [33] is an essential long-term financial management instrument. This involves optimizing the portfolio over multiple periods, taking into account not only the return and variance of each assessment but also the transaction costs. The objective function can be formulated as
$\max\left\{\sum_{t=1}^{T}\left(R_{t}-\frac{1}{2}\gamma\sigma_{t}^{2}-C_{t}\right)\right\},$ (8)
where $R_{t}$ is the return at time $t$, $\gamma$ is the risk aversion coefficient, $\sigma_{t}^{2}$ is the variance of the return, and $C_{t}$ represents the transaction costs.
Other Related Problems
In addition to valuation, mortality forecasting, and multi-period optimization, there are several other related problems of interest in the field of life insurance. First, underwriting risk [34] refers to the assessment of risk associated with insuring a new policyholder, and it often involves statistical and machine learning techniques to predict the likelihood of a claim. Additionally, longevity risk [35] is another crucial concern, as it represents the risk that policyholders live longer than expected, which can significantly impact the financial stability of life insurance companies. This issue is particularly relevant for annuity products. Furthermore, ensuring that life insurance products comply with regulatory standards is essential, as the standards can vary significantly across different jurisdictions. Finally, understanding and predicting customer behavior [36], such as lapse rates and claim frequencies, is critical for better managing and pricing insurance products.
III.3 Risk Management
Risk management is a critical aspect of financial and insurance industries, aiming to identify, assess, and prioritize risks followed by coordinated efforts to minimize, monitor, and control the probability or impact of unfortunate events. Effective risk management ensures the stability and profitability of financial institutions.
(conditional) Value at Risk
The Value at Risk (VaR) and conditional Value at Risk (cVaR) are common risk measures in risk management [37]. VaR is defined as the minimum loss bound at a given risk level $\alpha\in[0,1]$, such that the realized loss $x$ is less than this bound with a probability of at least $\alpha$. Mathematically, it is expressed as
$\text{VaR}_{\alpha}[X]=\inf{\{x|\mathbb{P}[X\leq x]\geq\alpha\}}.$ (9)
On the other hand, cVaR is the expected loss conditioned on the loss being under the bound $VaR_{\alpha}[X]$ defined as:
$\text{cVaR}_{\alpha}[X]=\mathbb{E}[X|0\leq X\leq\text{VaR}_{\alpha}[X]].$ (10)
Managing the total risk of an insurance portfolio is an essential problem. For each kind of insurance contract, one must decide whether to transfer the risk to the private market or to self-manage it within the insurance pool by allocating its reinsurance type.
Risk Transfer and Reinsurance
In the context of insurance, risk transfer and reinsurance are crucial strategies [7]. Risk transfer involves moving the risk from the insurance company to another party, typically through reinsurance contracts. Reinsurance helps insurers to manage their risk exposure and protect against significant losses. The decision on the type and extent of reinsurance involves evaluating the trade-offs between the cost of reinsurance and the benefits of reduced risk exposure.
Portfolio Risk Optimization
Portfolio risk optimization [38] is another important aspect of risk management. This involves optimizing the allocation of assets and liabilities to minimize risk while maximizing returns. Techniques such as mean-variance optimization and stochastic programming are commonly used. The objective function for portfolio optimization can be formulated as
$\max\left\{\sum_{i=1}^{n}\left(R_{i}-\frac{1}{2}\gamma\sigma_{i}^{2}\right)\right\},$ (11)
where $R_{i}$ is the return of asset $i$, $\gamma$ is the risk aversion coefficient, and $\sigma_{i}^{2}$ is the variance of the return of asset $i$.
Other Related Problems
There are several other related problems in risk management. First, credit risk management [39] involves assessing the risk of loss due to a borrower’s failure to make payments. This often includes the use of credit scoring models and default probability estimations. Second, operational risk management [40] focuses on risks arising from internal processes, systems, and external events, including fraud and cybersecurity threats. Furthermore, liquidity risk management [41] ensures that the institution can meet its short-term obligations without incurring significant losses.
IV Quantum algorithms for insurance
As aforementioned, modeling stochastic events is the core task of insurance and actuarial science. Quantum stochastic modeling involves embedding classical stochastic information into a quantum computer, processing the embedded data, and extracting the associated classical information for practical utilization. Efficiency in these steps is required for quantum advantages. Fortunately, quantum stochastic modeling could reduce the space exponentially with a further quadratic speedup. Firstly, statistical distributions, as well as initial states of stochastic processes, can be prepared in a quantum superposition, wherein the intrinsic symmetry of the underlying distribution can be employed to reduce the time complexity and mitigate the classical quantum input bottleneck. Secondly, for a stochastic process, the Markov dynamics described by a transition matrix can be simulated with quantum operations, herein the memory-less property can be employed to extend to continuous-time processes at a low cost. While the non-Markovian behaviors are related to an open quantum system, whose simulation has been widely studied [42, 43]. Finally, to evaluate the desired information from the expectation of moments, we implement the quantum-enhanced Monte-Carlo integration (QMC). It applies amplitude estimation to the quantum samplings described above with a quadratic accelerate [44, 45].
In spite of stochastic modeling, many practical insurance problems, such as insurance pricing and mortality forecasting, can be solved by data-driven methods of machine learning [32]. And quantum machine learning would be an appealing choice. In the NISQ era, quantum machine learning is realized via the well-known variational quantum algorithms [46], which are hybrid quantum-classical that combine parameterized quantum circuits as neural networks and classical computers as optimizers. Parameterized quantum circuits are set to be shallow to mitigate the hardware noise.
While in the FTQC era, where quantum circuits and the qubit number are greatly improved, it is possible to implement quantum machine learning algorithms based on linear algebra. The core subroutine in these algorithms is the solution of linear equations using the widely applied HHL algorithm [47, 48]. For instance, using quantum computers, The vision transformer [49], which is successfully applied on natural language processing tasks, can achieve speedups [50]. the singular value decomposition to extracting features can also be accelerated with quantum algorithms [51].
When considering life annuity and risk assessment management, optimization plays an essential role. Quadratic Unconstrained Binary Optimization (QUBO) problems can be translated into finding the ground state of Ising models, and a quantum annealer can derive the minima [52, 53, 54]. Variational optimization algorithms like the quantum approximation optimization algorithms [55] can also be applied for these goals.
V Results
In this part, we numerically and experimentally study several insurance problems. First, we develop a quantum algorithm to solve the excess evaluation problem. The quantum computing framework QPanda [56] is used for this task. Then we solve the reinsurance type allocation problem and the Lee-Carter model on “Wukong” [22], which is a high-performance superconducting quantum computer with the average relaxation time $T_{1}=14.84\mu s$ and $T_{2}=1.85\mu s$. The average fidelity of single-qubit gates and two-qubit gates are respectively $99.7\%$ and $96.3\%$.
V.1 Quantum excess evaluation
This part shows the numerical simulation results of the policy excess introduced before. We aim to model the payment for the reinsurer to pay $\mathbb{E}[\max(0,X-M)]$, or the discrete form:
$\frac{1}{N_{c}}\sum_{j=1}^{N_{c}}\max(0,X_{j}-M(X_{j})),$ (12)
where $N_{c}$ is the total claims. We assume $X_{j}$ to be the $j$-th loss followed the distribution
$f(x,\mu,\sigma)=\frac{1}{x\sigma\sqrt{2\pi}}e^{-\frac{(\ln x-\mu)^{2}}{2\sigma^{2}}}.$ (13)
$M(X_{j})$ is the payment, which is a non-decreasing function:
$M(x)=\begin{cases}0,&x\leq 1,\\ 0.6(x-1)+1,&x>1.\end{cases}$ (14)
As the policy number grows and the underlying distribution varies, a theoretical evaluation turns to be intractable and a numerical simulation demands a huge computation source consumption [2].
We developed a quantum algorithm for this problem, which is detailed in Sec. VII.1. In our simulation, we set $\mu=0$ and $\sigma=1$. Due to the convergence properties of the integral, it can be truncated with a certain level of precision. In this simulation, the excess $X_{j}$ is varied from 0 to 10 and The simulation results are shown in Fig. 1. As the number of qubits used to represent the distribution increases, the evaluated expectation becomes more accurate and convergent.
[…]
Figure 1: Simulation results of excess evaluation. The green and yellow lines represent the theoretical payment and its truncated version, respectively, demonstrating that truncating to $X\in[0,10]$ provides adequate accuracy. The blue dashed line shows the simulated results as a function of the number of qubits used. As the number of qubits increases, the simulation results converge towards greater accuracy.
V.2 Quantum reinsurance type allocation
As aforementioned, insurance companies pay reinsurance premiums for risk diversification and long-term stability with different risk characterizations and charges for the following two reinsurance types: Self-managed insurance pooling across different regions can reduce the insurance portfolio variance with lower premiums while being fragile to wide-ranged catastrophic [57, 58]. Private insurance purchasing from the international reinsurance market can further diversify risk at the cost of higher reinsurance premiums [59]. It is an essential yet intractable problem to allocate the reinsurance type for hundreds of insurance products to minimize the variance of the risk portfolio [60].
More formally, given the retained risk $R_{j}$ in the self-managed reinsurance pool that has not been transferred to the private insurance market, one aims to minimize the total variance of the total risk $R_{\text{total}}=\sum_{j}x_{j}R_{j}$ with varying reinsurance type $x_{j}\in\{0,1\}$ as
$\min_{\textbf{x}\in\{0,1\}^{n}}\text{Var}[R_{\text{total}}]\qquad\text{s.t.}\quad\sum_{j=1}^{n}x_{j}=pn.$ (15)
Herein the risk $R_{j}$ can be either a vector of historical data or a random variable from stochastic simulation, and the proportion $p$ satisfying $10<p<$ 1 represents the average allocation probability and is assumed to be fixed (for example, $p=0.5$ as in [60]).
We solve the case that the data is generated randomly. Noting that the loss function can be converted into an en expectation value of one Ising model through the standard transformation $x_{j}\rightarrow\frac{I-Z_{j}}{2}$ as
$H_{cost}=\sum_{j,k=1}^{n}v_{jk}\frac{I-Z_{j}}{2}\frac{I-Z_{k}}{2},$ (16)
[…]
We applied 6 superconducting qubits, illustrated in Fig. 2(a), for this task. To satisfy the constraint, instead of a traditional penalty function method, we utilize a more efficient problem-specified ansatz consisting of two parts: a parameterized initialization layer to prepare a quantum state satisfying the constraint, and repeating layers of reconfigurable Beam splitter (RBS) gates to transform the state while keeping the constraint unviolated [61, 62]. The RBS gate is $\text{diag}\left\{1,\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix},1\right\}$, which interacts between $|01\rangle$ and $|10\rangle$ while keeps $|00\rangle$ and $|11\rangle$ unchanged. A quantum circuit implementation of this gate is shown in Fig. 2(b). The number of RBS layers is set to be 3 and the whole ansatz after some simplifications is shown in Fig. 2(c).
[…]
(a)(b)(c)
Figure 2: (a) The 6 superconducting qubits used in the quantum computer “Wukong”. (b) The quantum circuit for the RBS gates. (c) The quantum circuit for the ansatz, with 6 qubits and 3 RBS layers, after some simplifications.
We apply a measurement-based error mitigation strategy to combat quantum hardware noise. While the theoretical output maintains the constraints, any deviation due to hardware noise is corrected by excluding erroneous shots from measurements.
The result is shown in Fig. 3. The bottom yellow line indicates the minimum orange of the loss function. We optimize the parameters based on the output of quantum computers while using the parameter to compute the noiseless loss function classically, which are labeled by “quantum_quantum” and “quantum_classical”, respectively. We can see that as the optimization goes, the loss function is decreased, and the corresponding “classical” loss function is closer to the target value, which is capable of generating the target $x$.
Figure 3: Experimental results for the reinsurance type allocation problem. The blue line illustrates the decrease in the loss function as outputted by quantum computers. The green line indicates the classically computed loss function using the same parameters. The orange line represents the target value.
V.3 Quantum Lee-Carter solver
Here, we report experimental results for solving the Lee-Carter model described in Eq. (7). Given a list of data $m_{x,t}$, the core task is to estimate $\beta_{x}$ and $\kappa_{t}$. This is usually done by first constructing the matrix $D$ whose elements are $D_{x,t}=\ln m_{x,t}-\alpha_{x}$. Then a singular value decomposition is performed on $D$, and the left(right) singular vector related to the maximum singular value is associated to $\beta_{x}(\kappa_{t})$. Then these data can be used for further predictions. We apply the quantum singular value decomposition algorithm [63]. A brief introduction of this algorithm is shown in Sec. VII.2.
We apply real-world data obtained from  [^1] https://www.mortality.org. The year $t$ is from $2014$ to $2017$ and the age $x$ is from 00 to $10-14$. Then the matrix size of $D$ is set as $4\times 4$. A total of 4 superconducting qubits are used and a total of 200 optimization steps are evaluated. The experimental results are shown in Fig. 4. It shows that the loss function decreases with the optimization steps. During this process, both the Frobenius norm distance and the Kullback-Leibler (K-L) divergence between the theoretical singular vector and our estimated singular vector also decrease, indicating the successful performance of the target singular value decomposition.
Figure 4: Experimental results of the Lee-Carter model. (a), (b), (c), and (d) respectively represent the evolution of the loss function, the Frobenius norm distance, and the K-L divergence (left and right) between the theoretical singular vector and our optimized singular vector over the epochs.
VI Conclusion and discussion
In the preceding sections, we delved into the realm of quantum computational insurance and actuarial science. We examined critical issues across the principal branches of insurance, including non-life insurance, life insurance, and reinsurance. Following this, we explored the application of quantum algorithms to address these challenges. Additionally, we conducted both numerical and experimental studies on several insurance problems using quantum algorithms. In this section, we outline the roadmap for the future of quantum computational insurance.
In the NISQ era, the available qubit number of quantum operation depth is limited due to the noise [20]. Although experiments have shown early quantum advantages on some specific applications [65], there is still a long way to universal quantum computing. While one of the most challenging problems is to strive for fault-tolerant quantum computing, how can the insurance industry make use of quantum computation in the NISQ era? We believe that both theoretical and experimental breakthroughs focusing on quantum insurance and actuarial science will be achieved. On the aspect of theoretical development, many quantum insurance algorithms will be proposed and proved to have quantum speed-up. Quantum stochastic modeling will be applied to study risk and ruin theory with many applications as suggested in the previous section. Also, quantum machine learning and optimization will be further developed on prediction and portfolio optimization problems. As for the aspect of experiments, quantum insurance algorithms, including but not limited to variational quantum algorithms and quantum Monte Carlo simulations, can be performed on quantum processors. The insurance industry can help to find the best scenarios and applications of quantum insurance and bring more attention to this cross field. With the cooperation of quantum computation scientists the insurance industry and academia, quantum algorithms driven or inspired by insurance can reveal more power of quantum computation and further applicability to insurance.
In the second stage of fault-tolerant quantum computing, errors on quantum processors can be corrected [66, 67]. These high-quality logical qubits, together with larger qubit numbers, will enable us to consider and perform more realistic quantum insurance applications [67]. As the power of quantum computation has been released, many exciting insurance applications and products can come true. The risk can be estimated and evaluated more precisely, and such an evaluation process can be even online and personalized, based on quantum computation with Internet and Internet of Things data. The potential loss can be predicted more early and precisely, benefiting from quantum computational climate predicting, mortality forecasting, and fat-tail events modeling, to name but a few. Furthermore, quantum natural language processing can help to accelerate complicated tasks such as AI customer service and consultation [68], and quantum pattern recognition can speed up claim analysis [69]. Nevertheless, one should also notice that, besides those quantum benefits mentioned above, there are quantum threats as well. Impacted by universal quantum computation, there will be lots of unexpected and consecutive influences that can be viewed as a source of uncertainty and risks. We believe that one of the most challenging problems is the modeling and evaluating of such a new source of risks. Once again, we caution that the insurance industry and quantum scientists should cooperate together to handle these challenges.
In this work, we have discussed the major ideas and impacts of quantum computational insurance and actuarial science, and have considered most of the possible applications and challenges at present and in the future. We also emphasized that we are far away from the final answer to those fundamental problems of quantum computational insurance and actuarial science. The applications and potential challenges within the context of quantum insurance should be invested in detail and considered carefully and the communication and cooperation of the insurance industry and quantum scientists are necessary.
Similar to other practical applications, quantum algorithms still face significant challenges. For instance, encoding a general distribution into a quantum computer requires the preparation of the corresponding quantum state, which demands exponential resources [70, 71]. Although $n$ qubits can encode $2^{n}$ amplitudes, each shot yields only a single $n$-bit binary string after performing the desired operations, thereby greatly limiting efficiency. Additionally, noise in quantum computers remains a challenge for most quantum algorithms. To address these issues, researchers are exploring several solutions. Quantum Random Access Memory [72] and advanced encoding strategies [73] can be utilized for efficient state preparation. Advanced sampling methods, such as shadow tomography [74, 75], offer faster readout. Furthermore, quantum error mitigation [76] strategies are being developed to combat noise in quantum hardware.
VII Methods
VII.1 Quantum excess evaluation algorithm
Here, we provide details about the quantum excess evaluation algorithm. The algorithm is divided into the following steps:
State preparation: When $n$ qubits are used to prepare the distribution of loss, we discrete the domain $[0,10]$ with $N=2^{n}$ points. Then the state is prepared as:
$|\psi_{1}\rangle=\sum_{j=0}^{N-1}\sqrt{p_{j}}|j\rangle,$ (17)
where $p_{j}=f(X_{j})/C$ with $C$ the normalization factor.
Quantum subtractor According to the definition of the payment function, we first need to evaluate $x-M(x)=0.4(x-1)$ when $x>1$. which needs to evaluate $x-1$ with a quantum Subtractor detailed in [77]. Since in the $n$-qubit case, the loss range 10 is divided into $12^{n}-$ 1 parts, then the threshold 1 can be represented by a n-bit binary representation as $1_{n}=\text{bin}(\frac{2^{n}-1}{10})$. With one auxiliary qubit, we can finish this task and the resulting state is:
$|\psi_{2}\rangle=\sum_{j=0}^{1_{n}-1}\sqrt{p_{j}}|N-j-1_{n}\rangle|1\rangle+\sum_{j=1_{n}}^{N-1}\sqrt{p_{j}}|j-1_{n}\rangle|0\rangle.$ (18)
Controlled rotations Apply a $X$ gate on the last qubit, and then add one new measurement qubit, followed by a $RY(\pi/2)$ gate, we have:
$\sum_{j=0}^{1_{n}-1}\sqrt{p_{j}}|N-x-1_{n}\rangle|0\rangle\left[\cos(\frac{\pi}{4})|0\rangle+\sin(\frac{\pi}{4})|1\rangle\right]$ (19)
$+$ $\sum_{j=1_{n}}^{N-1}|j-1_{n}\rangle|1\rangle\left[\cos(\frac{\pi}{4})|0\rangle+\sin(\frac{\pi}{4})|1\rangle\right].$
Then apply sequences of $CCRY$ gate, we get:
$\sum_{x=0}^{1_{n}-1}{\sqrt{p_{x}}|N-x-1_{n}\rangle|0\rangle}{(\cos\frac{\pi}{4}|0\rangle+\sin\frac{\pi}{4}|1\rangle)}$ (20)
$+$ $\sum_{x=1_{n}}^{N-1}{\sqrt{p_{x}}|x-1_{n}\rangle|1\rangle}\otimes(\cos\theta|0\rangle+\sin\theta|1\rangle),$
where $\theta=\frac{\pi}{4}+0.4c(x-1_{n})$ and $c$ is a small number.
Measure and post-processing Finally, measure the measurement qubit and the probability of obtaining $|1\rangle$ is
$\sum_{x=0}^{1_{n}-1}{p_{x}}{\sin^{2}\frac{\pi}{4}}+\sum_{x=1_{n}}^{N-1}{p_{x}}{\sin^{2}(\frac{\pi}{4}+0.4c(x-1_{n}))}$ (21)
$\approx$ $\frac{1}{2}+\sum_{x=1_{n}}^{N-1}{0.4c{p_{x}}(x-1_{n})}$
After some simple processing, we can get the desired result. The quantum circuit is shown in Fig. 5.
[…]
$\bm{\vdots}$
[…]
Figure 5: Quantum circuits for the quantum excess evaluation algorithm. The unitary operations $U_{X}$ and $U_{S}$ are used to prepare the state $|\psi_{1}\rangle$ and realize the subtraction operations. Followed controlling operations rotate the measurement qubit. The desired value is associated with the probability of obtaining $|1\rangle$.
VII.2 Quantum singular value decomposition algorithm
In this part, we briefly introduce the QSVD algorithm, which is detailed in [63]. The SVD of a matrix $D$ is $D=\sum_{i}\sigma_{i}|u_{i}\rangle\langle v_{i}|$, where $\sigma_{i}$ is the singular value, $|u_{i}\rangle/|v_{i}\rangle$ is the associated left/right singular vector. The QSVD algorithm starts with a vectorized representation of the matrix, which uses a vector of $MN$ elements to represent the $M\times N$ matrix.
$D$ $=\begin{pmatrix}D_{11}&\cdots&D_{1N}\\ \vdots&&\vdots\\ D_{M1}&\cdots&D_{MM}\end{pmatrix}$ (22)
$\to|D\rangle$ $=\begin{pmatrix}D_{11}&\cdots&D_{1N}&\cdots&D_{M1}&\cdots&D_{MM}\end{pmatrix}^{\operatorname{T}}.$
Therefore, the SVD of a matrix is transformed as
$D=\sum_{i}\sigma_{i}|u_{i}\rangle\langle v_{i}|\to|D\rangle=\sum_{i}\sigma_{i}|u_{i}\rangle|v_{i}\rangle.$ (23)
In quantum computers, with a re-normalization factor $C=\sqrt{\langle D|D\rangle}$ we can prepare the quantum state of $D$ in the Schmidt decomposition form
$|\psi\rangle_{AB}=\frac{1}{C}|D\rangle=\sum_{i}\lambda_{i}|u_{i}\rangle_{A}|v_{i}\rangle_{B},$ (24)
where $A,B$ represent the two-subsystems. Consider two unitary operations $U$ and $V$ that respectively convert the state $\{|u_{i}\rangle\}$ and $\{|v_{i}\rangle\}$ to the computational basis $\{|e_{i}\rangle\}$:
$U|u_{i}\rangle=e^{ip_{i}}|e_{i}\rangle,\qquad V|v_{i}\rangle=e^{iq_{i}}|e_{i}\rangle.$ (25)
We have
$(U_{A}\otimes V_{B})|\psi\rangle_{AB}=\sum_{i}\lambda_{i}e^{i(p_{i}+q_{i})}|e_{i}\rangle_{A}|e_{i}\rangle_{B}.$ (26)
Therefore, when measuring the two subsystems $A$ and $B$ of the state $|\psi\rangle_{AB}$, we will get the same measurement result in the computational basis. The corresponding probability is $\lambda_{i}^{2}=\sigma_{i}^{2}/C^{2}$. Finally, apply $U^{\dagger}$ and $V^{\dagger}$ on the computational basis state $|e_{i}\rangle$ will result in the singular vector $|u_{i}\rangle$ and $|v_{i}\rangle$.
To obtain $U/V$. we apply the variational QSVD algorithm by introducing an optimization procedure. That is: after preparing the state $|\psi\rangle_{AB}$, we apply two parameterized quantum circuits $U_{A}(\bm{\theta}_{1})$ and $V_{B}(\bm{\theta}_{2})$. The loss function is
$C=\frac{1}{N_{m}}\sum_{x=1}^{N_{m}}\delta(p_{x},q_{x}),$ (27)
where $N_{m}$ is the number of shots, $p_{x},q_{x}$ are the two binary strings at the $x$-th shot. $\delta(a,b)$ is 0 if $a\neq b$ and 1 if $a=b$. A classical optimizer is employed to optimize the parameters to minimize the loss function.
ACKNOWLEDGEMENT
This work has been supported by the National Key Research and Development Program of China (Grant No. 2023YFB4502500), the National Natural Science Foundation of China (Grant No. 12404564), and the Aeronautical Science Foundation of China (Grant No. 2022Z073004001).